OpenStudy (anonymous):

Find the volume of a pyramid with an area of 15 and a height of 10.

OpenStudy (wolf1728):

I'm assuming it is a 4 sided pyramid and the area is the area of all four sides.

OpenStudy (wolf1728):

The area of 1 side is 3.75

OpenStudy (wolf1728):

Base length would be .7495 Volume = (base * height) / 3 Volume = (base²*height)/3 = .7495*.7495*10/3 = 1.8725